COMPUTE IN MATLAB ONLY!!!!!!

  1. Compute the total charge contained in a cylindrical tube oriented along the z-axis. The line charge density is rl = 2*z, where z is the distance in meters from the bottom end of the tube. The tube length is 10 cm.
Answer #1

So We have Length charge density as =2 *z C/m

We know that charge contained in a length charge density is given as

Here a=0 b=10cm=0.1m

  

  

  MATLAB CODE

 syms z Rhoz = 2*z; Q= int(Rhoz,[0 0.1]); fprintf('Total charge contained in cylinder is : %0.3f Columbs \n',Q);

RESULT

Total charge conatined in cylinder is : 0.010 Columbs
>>